# Online Linear Quadratic Tracking with Regret Guarantees

### v2

@@ 2 Problem Statement
There exist $\bar{R}\in\mathbb{R}_{+}$, such that $\|r_{t}\|\leq\bar{R}$ for all $t\in\mathbb{N}$.
@@ 4 Regret Analysis
where $u_{t}^{\pi_{2}}$ is the input generated by the policy $\pi_{2}$ at time t, $e_{t}^{\pi_{2}}$ is the state at time $t$ generated by applying the policy $\pi_{2}$, and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{Q}}_{t}^{\pi_{1}}(e,u)=\|e\|_{Q}^{2}+\|u\|_{R}^{2}+J_{t+1}(Ae+Bu+w_{t},u^{\pi_{1}})$ is the Q-function for policy $\pi_{1}$ and $J_{i}(e_{i},u)$ is the cost-to-go at time step $i$, with initial state $e_{i}$ and control signal $u$.
@@
$\mathcal{R}^{\pi}(w,e_{0})=\sum_{t=0}^{T-1}\left(u_{t}^{\pi}-u_{t}^{*}\right)^{\top}\left(R+B^{T}PB\right)\left(u_{t}^{\pi}-u_{t}^{*}\right).$
@@
${\mathcal{R}^{\pi}(w,e_{0})=\sum_{t=0}^{T-1}\mathcal{Q}_{t}^{*}(e_{t}^{\pi},u_{t}^{\pi})-\min_{u}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{Q}}_{t}^{*}(e_{t}^{\pi},u),}$ (16)
@@
where the last equality follows from the closed form of $J_{t+1}(x,u^{\star})$ as an extended quadratic function of $x$ [12, 17]. Thus, since $u_{t}^{\star}$ minimizes an extended quadratic function, $\mathcal{Q}_{t}^{*}(e_{t}^{\pi},u_{t}^{\pi})-\mathcal{Q}_{t}^{*}(e_{t}^{\pi},u_{t}^{\star})=\|u_{t}^{\pi}-u_{t}^{*}\|_{\left(R+B^{T}PB\right)}$.
@@
$\left\|u^{\mathrm{ogd}}_{t}-u_{t}^{*}\right\|=\left\|-Ke_{t}+v_{t}+Ke_{t}+\sum_{i=t}^{T-1}K_{w}^{i,t}w_{i}\right\|$

### v3

@@ 2 Problem Statement
There exists a $\bar{R}\in\mathbb{R}_{+}$, such that $\|r_{t}\|\leq\bar{R}$ for all $t\in\mathbb{N}$.
@@ 4 Regret Analysis
where $e_{t}^{\pi_{2}}$ is the state at time $t$ achieved by applying the policy $\pi_{2}$, $u_{t}^{\pi_{2}}$ is the input generated by the policy $\pi_{2}$ at time t, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{Q}}_{t}^{\pi_{1}}(e,u)=\|e\|_{Q}^{2}+\|u\|_{R}^{2}+J_{t+1}(Ae+Bu+w_{t},u^{\pi_{1}})$ is the Q-function for policy $\pi_{1}$ and $J_{i}(e_{i},u)$ is the cost-to-go at time step $i$, with initial state $e_{i}$ and control signal $u$.
@@
$\mathcal{R}^{\pi}(w,e_{0})=\sum_{t=0}^{T-1}\left(u_{t}^{\pi}-u_{t}^{\star}\right)^{\top}\left(R+B^{T}PB\right)\left(u_{t}^{\pi}-u_{t}^{\star}\right),$
where $u_{t}^{\pi}$ and $u_{t}^{\star}$ denote the inputs generated by $\pi$, and the optimal policy, both evaluated at the policy state $e_{t}^{\pi}$ at time $t$.
@@
${\mathcal{R}^{\pi}(w,e_{0})=\sum_{t=0}^{T-1}\mathcal{Q}_{t}^{\star}(e_{t}^{\pi},u_{t}^{\pi})-\min_{u}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{Q}}_{t}^{\star}(e_{t}^{\pi},u),}$ (16)
@@
where the last equality follows from the closed form of $J_{t+1}(x,u^{\star})$ as an extended quadratic function of $x$ [12, 17]. Thus, since $u_{t}^{\star}$ minimizes an extended quadratic function, $\mathcal{Q}_{t}^{\star}(e_{t}^{\pi},u_{t}^{\pi})-\mathcal{Q}_{t}^{\star}(e_{t}^{\pi},u_{t}^{\star})=\|u_{t}^{\pi}-u_{t}^{\star}\|^{2}_{\left(R+B^{T}PB\right)}$.
@@
$\left\|u^{\mathrm{ogd}}_{t}-u_{t}^{\star}\right\|=\left\|-Ke_{t}+v_{t}+Ke_{t}+\sum_{i=t}^{T-1}K_{w}^{i,t}w_{i}\right\|$